OpenStudy (freemap):

Which of the following best describes the thermosphere? A.The thermosphere sits directly below the troposphere. B. The atmosphere is so thin that gas molecules rarely contact each other. C. Lighter gases such as nitrogen and oxygen rise to the top of the thermosphere. D. Heavier gases such as hydrogen and helium stay at the bottom of the thermosphere.

OpenStudy (freemap):

@Michele_Laino

OpenStudy (michele_laino):

I am sorry, I am not good with the study of thermosphere

OpenStudy (freemap):

ok, thanks for viewing, appreciate it. I think its C or B though, anyway thanks

OpenStudy (michele_laino):

ok! :)

OpenStudy (freemap):

Would you happen to know this ? A car starts from rest and accelerates uniformly over a time of 5.21 seconds for a distance of 110 m. Determine the acceleration of the car.

OpenStudy (michele_laino):

here we have to apply this formula: \[s = \frac{1}{2}a{t^2}\] hence: \[a = \frac{{2s}}{{{t^2}}} = \frac{{2 \cdot 110}}{{{{5.21}^2}}} = ...?\]

OpenStudy (freemap):

8.10

OpenStudy (michele_laino):

that's right!

OpenStudy (freemap):

Thank you

OpenStudy (michele_laino):

:)